OpenStudy (anonymous):

Solve, using the substitution method. 4x + 2y = 11 x – 2 = –2y Help?

OpenStudy (igreen):

First, rearrange the second equation: x - 2 = -2y Add 2 to both sides: x = -2y + 2 Okay, now we can plug in -2y + 2 for x in the first equation: 4x + 2y = 11 4(-2y + 2) + 2y = 11 -8y + 8 + 2y = 11 -6y + 8 = 11 -6y = 3 Can you divide 3 / -6?